fractional equations and extraneous solutions
2 / x + 5 - 1 / x - 3 = 3

Answer by checkley79(3341)

(2 / x) + 5 - (1 / x) - 3 = 3
2/X-1/X=3-5+3
1/X=1
X=1 ANS.
OR:
2/(X+5)-(1/X-3)=3
[2(X-3)-1(X+5)]/(X+5)(X-3)]=3
[(2X-6-X-5)/(X^2+2X-15)]=3
[(3X-11)/(X^2+2X-15)]=3 CROSS MULTIPLY
3(X^2+2X-15)=3X-11
3X^2+6X-45-3X+11=0
3X^2+3X-45+11=0
4X^2+3X-34=0
SOL. (-3.314 & 2.564)